# Exploring Transferability of Multimodal Adversarial Samples for Vision-Language Pre-training Models with Contrastive Learning

### v3

@@ I Introduction
Vision-Language Pre-training (VLP) models are essential for integrating visual and textual data, which is considered as a significant focus of current research aimed at bridging the gap between vision and language understanding [1, 2, 3, 4]. The contrastive learning effectively aligns cross-modal feature representations by attracting matched image-text pairs and repelling non-matched pairs, demonstrating a profound capacity to learn generalizable and transferable multimodal features without reliance on labeled data [5]. The efficacy of this alignment strategy in VLP models has led to substantial advancements in various applications [6, 7, 5, 1, 2], notably in image-text retrieval, visual entailment, visual question answering, and visual grounding, showcasing their transformative potential in multimodal understanding.
@@
It is well studied that adversarial robustness can introduce vulnerabilities through minor adversarial perturbations to inputs, whether textual [10, 9], visual [11, 12, 13, 14] or audio [15]. Based on it, transfer-based adversarial attacks are proposed which are considered more practical in real-world applications [8, 16, 17]. Some preliminary efforts have been undertaken to explore the adversarial robustness of the VLP models. The co-attack method proposed to give a sequential text-then-image approach to generate multi-modal adversarial samples in a white-box setting [18]. The robustness of the large VLP models is evaluated under the task of image-text generation and considers the transferability in a box-box style. Though comprehensive, this evaluation study is limited to the single-modal adversarial sample, namely image inputs [19]. SGA [20] extends the Co-attack framework by incorporating data augmentation with cross-modal guidance. This enhancement facilitates the transferability of multimodal adversarial samples through sequential optimization.
It is worth noting that currently the adversarial attacks in VLP models lack the intricate interplay between text and image modalities. For effective modality alignments, multimodal adversarial attacks must target each modality individually and meanwhile preserve the contextual integrity of the adversarial pair, as shown in Figure 1. Due to the discrete nature of texts and contrasts sharply with the continuous nature of images, developing a unified framework for generating the transferable adversarial text-image pair represents an unexplored yet critical area of research.
In this paper, we propose a gradient-based multimodal adversarial attack method with contrastive learning (VLP-attack) to generate transferable multimodal adversarial samples against VLP models. Specifically, we combine the textual adversarial attack and visual adversarial attack into a unified framework based on gradients, which can produce the adversarial texts and the adversarial images simultaneously by gradients under perceptibility constraints. Moreover, we use image-text contrastive loss and intra-modal contrastive loss to improve the transferability of multimodal adversarial samples. We apply the general paradigm in VLP models: CLIP as the surrogate model and prove our method outperforms the existing single-modal transfer-based adversarial attack methods [9, 21, 22, 23, 19] and multimodal adversarial attack method [18, 20] in the image-text retrieval task and the visual entailment task where the victim models include CLIP series [24, 25], ALBEF [1], TCL [2], BLIP [3], BLIP2 [4], and MiniGPT-4 [26].
@@ II-B Black-box Adversarial Attack and Transferability
Black-box adversarial attacks and their transferability have gained significant attention in recent years due to the vulnerability of machine-learning models to adversarial samples [31, 32, 33, 34]. Compared with the white-box adversarial attack [11, 35], the black-box attack assuming no knowledge about the target model is a harder setting to study adversarial transferability. In the transfer-based black-box setting, the text adversarial attack methods: Bert-attack [9], PWWS [36], GBDA [21], and image adversarial attack methods: MI-FGSM [22], TI-FGSM [8], SI-NI-FGSM [23], C-GSP [37] and [38, 39], generate adversarial samples on surrogate models at a white-box setting to attack other target models, even when the attacker has no access to the target model’s architecture, gradients, or parameters. However, given the complexity of multimodal data, traditional transfer-based strategies against vision-only models, designed for single-modal data, fall short in effectively perturbing the alignment between text and image features in VLP models. Crafting multimodal adversarial samples with high transferability thus presents a significant challenge.
@@ III-A Problem Formulation
Given an image-text pair, our goal is to minimize the semantic similarity between the image and the corresponding text with minor adversarial perturbations on images and texts on surrogate models. Let $x_{i}$ denote an original image, $x_{t}$ denote an original text, $x^{\prime}_{i}$ denote an adversarial image, $x^{\prime}_{t}$ denote an adversarial text, and $y^{t}$ denote the corresponding ground-truth label in the image-text retrieval task and the visual entailment task. We use $\mathcal{F}_{s}$ to denote the surrogate VLP models and $\mathcal{F}_{v}$ to denote the victim VLP models, where the output of $\mathcal{F}_{s}$ is the feature representation of the input and the output of $\mathcal{F}_{v}$ is the prediction score. To generate valid multimodal adversarial samples, we need to minimize the cosine similarity Loss function $J$:
$L_{adv}=\left\{\begin{array}[]{cl}\mathrm{min}J(\mathcal{F}_{s}(x^{\prime}_{i}),\mathcal{F}_{s}(x^{\prime}_{t}))\\ s.t.||x^{\prime}_{i}-x_{i}||_{\infty}\leq\epsilon\\ s.t.\mathrm{similarity}(x^{\prime}_{t},x_{t})\leq\beta\\ \end{array}\right.$ (1)
on surrogate models and make $\mathcal{F}_{v}((x^{\prime}_{i},x^{\prime}_{t}))=y^{adv}$ ($y^{t}\neq y^{adv}$) on victim models. For the image $x^{\prime}_{i}$, under the constraint, the adding perturbation should look visually similar to the benign image $x_{i}$. For the text $x^{\prime}_{t}$, we modify the words in a benign text $x_{t}$, which can not affect the original semantic understanding for humans. In this work, we use the $L_{\infty}$ norm to constrain the perceptibility of adversarial perturbations, i.e., $||x^{\prime}_{i}-x_{i}||_{\infty}\leq\epsilon$. For the fluency of $x^{\prime}_{t}$ , we constrain the $x^{\prime}_{t}$ with the objective of the next token prediction by maximizing the likelihood given previous tokens. For the semantic similarity between $x_{t}$ and $x^{\prime}_{t}$, we use the bert score to constrain the adversarial text $x^{\prime}_{t}$. The framework of VLP-attack is shown in Figure 2.
@@ III-B The objective for multimodal Adversarial Perturbation
The optimization objective of multimodal adversarial perturbation consists of three components: image-text semantic similarity loss $L_{adv}$ for multimodal adversarial sample, soft constraints for adversarial texts, and contrastive loss for transferability, where $L_{adv}$ has been introduced in section 3.1.
@@ III-B1 Soft Constraints for Adversarial Texts
Adversarial Text Distribution. Due to the discrete nature of text data, the first step of optimizing both the adversarial texts and the adversarial images simultaneously based on gradients is to transform the text into a continuous distribution. Motivated by [21], we can obtain the text distribution through the Gumble-softmax [40]. Specifically, for a sequence of words $x_{t}=[w_{1},w_{2},...w_{n}]$ where each $w_{j}\in\mathcal{V}$ and $\mathcal{V}$ is a fixed vocabulary, we use a Gumble-softmax distribution $P_{\theta}$ parameterized by a matrix $\Theta\in R^{n\times V}$ to sample a sequence of probability vectors $\pi$, where $\pi=\pi_{1}...\pi_{n}$, and use an embedding function $e(\cdot)$ to convert $\pi$ to input text embedding. Samples $\pi$ and adversarial text features $e(\pi)$ are obtained according to the process:
$(\pi_{k})_{j}:=\frac{\exp((\Theta_{k,j}+g_{k,j})/t)}{\sum_{v=1}^{V}\exp((\Theta_{k,v}+g_{k,v})/t)}$ (2)
@@
where $g_{k,j}\sim$Gumble (0,1) and $t>0$ is a temperature parameter that controls the smoothness of Gumble-softmax distribution.
@@
where $\mathrm{log}p_{dis}$ is the cross-entropy between the next word distribution and the previously predicted word distribution. Meanwhile, for the semantic similarity between adversarial texts and origin texts, we use BERT score [41] to constraint the semantic gap:
$L_{sim}(x_{t},x^{\prime}_{t})=\sum_{k=1}^{n}w_{k}\underset{j=1,...,m}{\mathbf{max}}\phi(x_{t})_{k}^{\top}\phi(x^{\prime}_{t})_{j}$ (5)
where $x_{t}=x_{t,1}\cdot\cdot\cdot x_{t,n}$ and $x^{\prime}_{t}=x^{\prime}_{t,1}\cdot\cdot\cdot x^{\prime}_{t,m}$. $\phi(\cdot)$ is a language model that can generate contextualized embeddings for the token sequences.
@@ III-B2 Contrastive Loss for Transferability
$L_{nce}(x_{i},x^{+}_{t},x^{-}_{t})=E\left[\log\frac{e^{(sim(x_{i},x^{+}_{t})/\tau)}}{\sum^{K}_{k=1}{e^{(sim(x_{i},\hat{x}_{t,k})/\tau)}}}\right]$ (6)
where $\tau$ is a temperature hyper-parameter, $\hat{x}_{t}$ is a set of all positive samples and negative samples for benign image $x_{i}$.
Cross-modal contrastive Loss. In image-text contrastive learning, our goal is to minimize the semantic similarity between the adversarial image and adversarial text that is matched in the original space. In image-to-text loss, the original text and text augmentations are taken as negative samples $x^{-}_{t}$. In text-to-image loss, the original image and image augmentations are regarded as negative samples $x^{-}_{i}$. Intuitively, we encourage the features of the adversarial images and benign textual features to be unaligned in the embedding space, which can affect the multimodal features fusion in turn. Taken the image-to-text contrastive loss and text-to-image contrastive loss together as follows:
$L_{itm}=\frac{1}{2}\left[L_{nce}(x^{\prime}_{t},x^{+}_{i},x^{-}_{t})+L_{nce}(x^{\prime}_{i},x^{+}_{t},x^{-}_{i})\right]$ (7)
@@
Intra-modal Contrastive Loss. we utilize intra-modal contrastive learning to push adversarial samples semantically different from the benign samples in the same modality. Specifically, we consider some random views ($x_{i,1},x_{i,2},...,x_{i,n}$) of the benign image $x_{i}$ under random data augmentation as negative samples $x^{-}_{i}$, and randomly sample images from the test set as the positive samples $x^{+}_{i}$.
$L_{i2i}=L_{nce}(x^{\prime}_{i},x^{+}_{i},x^{-}_{i})$ (8)
@@
where $a,b,c,d,g$ are the hyper-parameters that can control the strength of different constraint terms. Our purpose is to maximize the $L(\Theta,x^{\prime}_{i})$ by optimizing the matrix $\Theta$ for adversarial text $x^{\prime}_{t}$ and the perturbation on the image $x^{\prime}_{i}$.
@@ III-C The generation of multimodal adversarial samples
$\left\{\begin{array}[]{cl}g^{j}_{\Theta}=\nabla_{\Theta}L(\mathcal{F}_{s}(x^{j}_{i}),e(\pi^{j})))\\ \Theta^{j+1}=f(\Theta^{j},\alpha_{\Theta},g^{j}_{\Theta},m^{j}_{\Theta})\end{array}\right.$ (10)
@@
For the images that are continuous data, similar to the momentum updating approach in [23], we use a variation of normal gradient descent to make a jump in the direction of previously accumulated gradients before computing the gradients in each iteration. Meanwhile, following the [16], we diverse the pattern of $x_{i}^{j}$. we input ($x^{\prime}_{i},x^{\prime}_{t}$) to $\mathcal{F}_{s}$ and obtain the gradient of $x^{\prime}_{i}$ :
$\left\{\begin{array}[]{cl}x_{i}^{nes}=T(x^{j}_{i},p)+g_{x^{\prime}_{i}}^{j}\\ g^{j+1}_{x^{\prime}_{i}}=\mu\cdot g^{j}_{x^{\prime}_{i}}+\frac{\nabla_{x^{\prime}_{i}}L(\mathcal{F}_{s}(x^{nes}_{i}),e(\pi^{j})))}{||\nabla_{x^{\prime}_{i}}L(\mathcal{F}_{s}(x^{nes}_{i}),e(\pi^{j}))||_{1}}\\ x^{j+1}_{i}=\mathrm{Clip}^{\epsilon}_{x_{i}}\{x^{j}_{i}+\alpha_{x^{\prime}_{i}}\cdot\mathrm{sign}(g^{j+1}_{x^{\prime}_{i}})\}\\ \end{array}\right.$ (11)
where $T$ is the stochastic transformation function with the probability $p$, $\alpha_{x^{\prime}_{i}}$ is the step size, $\mu$ is the decay factor of the momentum term, $g^{j}_{x^{\prime}_{i}}$ is the accumulated gradient at iteration $j$, and $\mathrm{Clip}^{\epsilon}_{x_{i}}$ indicates the resulting adversarial images are clipped within the $\epsilon-$ball of the original image $x_{i}$.
@@ IV-A Experimental Settings
Evaluation Metrics. For evaluation, we report (1) the attack success rate (ASR), which measures the rate of successful adversarial samples. Similar to Co-attack [18], we define the ASR as the results when the original samples are fed into the target models minus the results when the adversarial samples are fed into the target models. (2) Semantic Similarity (Sim.), which is computed between the original and adversarial sentences and is commonly approximated by the universal sentence encoder [45]. (3) Tokens error rate (TER.), which is the proportion of words or characters in a text that have been altered to create an adversarial text, generally, less perturbation results in more semantic consistency. The whole method is implemented by Pytorch [46] and all experiments are conducted in a GeForce RTX A6000 GPU.

### v5

@@ I Introduction
Vision-Language Pre-training (VLP) models are pivotal in integrating visual and textual data, which is considered as a significant research focus aimed at enhancing vision-language understanding [1, 2, 3, 4]. Contrastive learning aligns cross-modal features effectively by attracting matched image-text pairs and repelling mismatched ones, demonstrating its capability to learn transferable multimodal features without labeled data [5]. This alignment strategy in VLP models has led to substantial advancements in applications such as image-text retrieval [6], visual entailment [7], and visual question answering [8], highlighting their transformative potential in multimodal understanding.
@@
It is well studied that the adversarial robustness would bring vulnerability by adding the minor input adversarial perturbations, either textual [11, 10], or visual [12, 13]. Based on it, transfer-based adversarial attacks are proposed which are considered more practical in real-world applications [9, 14, 15, 16]. Some preliminary efforts have been undertaken to explore the adversarial robustness of the VLP models. The Co-attack method proposed to give a sequential text-then-image approach to generate multi-modal adversarial samples in a white-box setting [17]. The SGA approach extends the Co-attack framework by leveraging cross-modal interactions across multiple alignments with set-level guidance, enhancing the transferability of multimodal adversarial samples in the black-box setting [18]. The robustness of the large VLP models is evaluated under the task of image-text generation and considers the transferability in a box-box style. Though comprehensive, this evaluation study is limited to the adversarial images [19].
It is worth noting that currently adversarial strategies in VLP models often overlook the complex interplay between text and image modalities. Effective multimodal adversarial attacks must target each modality while considering the contextual relation of the adversarial pair, which enhances their efficacy in VLP models, as depicted in Figure 1. Given the discrete nature of texts and the continuous nature of images, developing a unified framework for generating transferable adversarial text-image pairs remains a critical and unexplored area.
In this paper, we propose a gradient-based multimodal adversarial attack method with contrastive learning (VLP-attack) to generate transferable multimodal adversarial samples against VLP models. Our approach unifies textual and visual adversarial attacks within a gradient-based framework, enabling simultaneous production of adversarial texts and images under perceptibility constraints. We employ image-text and intra-modal contrastive losses to enhance sample transferability. Our method is validated using CLIP as the surrogate model, outperforming existing single-modal and multimodal adversarial attack methods in image-text retrieval and visual entailment tasks. Victim models tested include the CLIP series [20, 21], ALBEF [1], TCL [2], BLIP [3], BLIP2 [4], and MiniGPT-4 [22].
@@ II-B Black-box Adversarial Attack and Transferability
Black-box adversarial attacks and their transferability have gained significant attention in recent years due to the vulnerability of machine-learning models to adversarial samples [28, 29, 30, 31]. Compared with the white-box adversarial attack [12, 32], the black-box attack assuming no knowledge about the target model is a harder setting to study adversarial transferability. In the transfer-based black-box setting, the text adversarial attack methods: Bert-attack [10], PWWS [33], GBDA [34], and image adversarial attack methods: MI-FGSM [35], TI-FGSM [9], SI-NI-FGSM [26], C-GSP [36] and [37, 38], generate adversarial samples on surrogate models at a white-box setting to attack other target models, even when the attacker has no access to the target model’s architecture, gradients, or parameters. Among them, LPM [39] posits that model-specific discriminative regions contribute significantly to overfitting on the source model. To address this, it employs a differential evolutionary algorithm to identify a patch-wise binary mask. This mask is used to drop out image patches associated with these model-specific regions, thereby enhancing the transferability of adversarial examples. However, given the complexity of multimodal data, traditional transfer-based strategies against vision-only models, designed for single-modal data, fall short in effectively perturbing the alignment between text and image features in VLP models. Crafting multimodal adversarial samples with high transferability thus presents a significant challenge.
@@ III-A Problem Formulation
Given an image-text pair, our goal is to minimize the semantic similarity between the image and the corresponding text with minor adversarial perturbations on images and texts on surrogate models. Let $i$ denote an original image, $t$ denote an original text, $i^{\prime}$ denote an adversarial image, $t^{\prime}$ denote an adversarial text, and $y$ denote the corresponding ground-truth label in the image-text retrieval task and the visual entailment task. We use $\mathcal{F}_{s}$ to denote the surrogate VLP models and $\mathcal{F}_{v}$ to denote the victim VLP models, where the output of $\mathcal{F}_{s}$ is the feature representation of the input and the output of $\mathcal{F}_{v}$ is the prediction score. To generate valid multimodal adversarial samples, we need to minimize the cosine similarity Loss function $J$:
$L_{adv}=\left\{\begin{array}[]{cl}\mathrm{min}J(\mathcal{F}_{s}(i^{\prime}),\mathcal{F}_{s}(t^{\prime}))\\ s.t.||i^{\prime}-i||_{\infty}\leq\epsilon\\ s.t.\mathrm{similarity}(t^{\prime},t)\leq\beta\\ \end{array}\right.$ (1)
on surrogate models and make $\mathcal{F}_{v}((i^{\prime},t^{\prime}))=y^{adv}$ ($y\neq y^{adv}$) on victim models. For the image $i^{\prime}$, under the constraint, the adding perturbation should look visually similar to the benign image $i$. For the text $t^{\prime}$, we modify the words in a benign text $t$, which can not affect the original semantic understanding for humans. In this work, we use the $L_{\infty}$ norm to constrain the perceptibility of adversarial perturbations, i.e., $||i^{\prime}-i||_{\infty}\leq\epsilon$. For the fluency of $t^{\prime}$ , we constrain the $t^{\prime}$ with the objective of the next token prediction by maximizing the likelihood given previous tokens. For the semantic similarity between $t$ and $t^{\prime}$, we use the bert score to constrain the adversarial text $t^{\prime}$. The framework of VLP-attack is shown in Figure 2.
@@ III-B The objective for multimodal Adversarial Perturbation
The optimization objective of multimodal adversarial perturbation consists of three components: image-text semantic similarity loss $L_{adv}$ for multimodal adversarial sample, soft constraints for adversarial texts, and contrastive loss for transferability, where $L_{adv}$ has been introduced in section III-A.
@@ III-B1 Soft Constraints for Adversarial Texts
Adversarial Text Distribution. Due to the discrete nature of text data, optimizing adversarial texts and images simultaneously requires transforming the text into a continuous distribution. Following the approach in Guo et al.[34], we employ the Gumbel-Softmax distribution[40] to model text data. Specifically, for a sequence of words $t=[w_{1},w_{2},...,w_{n}]$ with each $w_{j}$ belonging to a fixed vocabulary $\mathcal{V}$, we utilize a Gumbel-Softmax distribution $P_{\theta}$, parameterized by a matrix $\Theta\in\mathbb{R}^{n\times V}$. This distribution is used to sample a sequence of probability vectors $\pi$, where $\pi=\pi_{1}...\pi_{n}$ ($\pi_{n}$ denotes the probability vector of $w_{n}$ ). These vectors are then converted into input text embeddings using an embedding function $e(\cdot)$. Probability vectors $\pi$ and adversarial text features $e(\pi)$ are obtained according to the process:
$(\pi_{k})_{j}:=\frac{\exp((\Theta_{k,j}+g_{k,j})/\tau)}{\sum_{v=1}^{V}\exp((\Theta_{k,v}+g_{k,v})/\tau)}$ (2)
@@
where $g_{k,j}\sim$Gumble (0,1) and $\tau>0$ is a temperature parameter that controls the smoothness of Gumble-softmax distribution.
@@
where $\mathrm{log}p_{dis}$ is the cross-entropy between the next word distribution and the previously predicted word distribution. Meanwhile, to ensure that modifications in adversarial texts do not alter the original text’s semantic interpretation for humans, we employ the BERT score [41] to maintain semantic consistency and constrain the semantic gap:
$L_{sim}(t,t^{\prime})=\sum_{k=1}^{n}w_{k}\underset{j=1,...,m}{\mathbf{max}}\phi(t)_{k}^{\top}\phi(t^{\prime})_{j}$ (5)
where $t=t_{1}\cdot\cdot\cdot t_{n}$ and $t^{\prime}=t^{\prime}_{1}\cdot\cdot\cdot t^{\prime}_{m}$. $\phi(\cdot)$ is a language model that can generate contextualized embeddings for the token sequences.
@@ III-B2 Contrastive Loss for Transferability
$L_{nce}(i,t^{+},t^{-})=E\left[\log\frac{e^{(sim(i,t^{+})/\tau)}}{\sum^{K}_{k=1}{e^{(sim(i,\hat{t}_{k})/\tau)}}}\right]$ (6)
where $\tau$ is a temperature hyper-parameter, $\hat{t}$ is a set of all positive samples and negative samples for benign image $i$.
Cross-modal contrastive Loss. In image-text contrastive learning, our goal is to minimize the semantic similarity between the adversarial image and adversarial text that is matched in the original space. In image-to-text loss, the original text and text augmentations are taken as negative samples $t^{-}$. In text-to-image loss, the original image and image augmentations are regarded as negative samples $i^{-}$. Intuitively, we encourage the features of the adversarial images and benign textual features to be unaligned in the embedding space, which can affect the multimodal features fusion in turn. Taken the image-to-text contrastive loss and text-to-image contrastive loss together as follows:
$L_{itm}=\frac{1}{2}\left[L_{nce}(t^{\prime},i^{+},t^{-})+L_{nce}(i^{\prime},t^{+},i^{-})\right]$ (7)
@@
Intra-modal Contrastive Loss. we utilize intra-modal contrastive learning to push adversarial samples semantically different from the benign samples in the same modality. Specifically, we consider some random views ($i_{1},i_{2},...,i_{n}$) of the benign image $i$ under random data augmentation as negative samples $i^{-}$, and randomly sample images from the test set as the positive samples $i^{+}$.
$L_{i2i}=L_{nce}(i^{\prime},i^{+},i^{-})$ (8)
@@
where $a,b,c,d,g$ are the hyper-parameters that can control the strength of different constraint terms. Our purpose is to maximize the $L(\Theta,i^{\prime})$ by optimizing the matrix $\Theta$ for adversarial text $t^{\prime}$ and the perturbation on the image $i^{\prime}$.
@@ III-C The generation of multimodal adversarial samples
$\left\{\begin{array}[]{cl}g^{j}_{\Theta}=\nabla_{\Theta}L(\mathcal{F}_{s}(i^{j}),e(\pi^{j})))\\ \Theta^{j+1}=f(\Theta^{j},\alpha_{\Theta},g^{j}_{\Theta},m^{j}_{\Theta})\end{array}\right.$ (10)
@@
For the images that are continuous data, similar to the momentum updating approach in [26], we use a variation of normal gradient descent to make a jump in the direction of previously accumulated gradients before computing the gradients in each iteration. Meanwhile, following the [14], we diverse the pattern of $i^{j}$. we input ($i^{\prime},t^{\prime}$) to $\mathcal{F}_{s}$ and obtain the gradient of $i^{\prime}$ :
$\left\{\begin{array}[]{cl}x_{i}^{nes}=T(i^{j},p)+g_{i^{\prime}}^{j}\\ g^{j+1}_{i^{\prime}}=\mu\cdot g^{j}_{i^{\prime}}+\frac{\nabla_{i^{\prime}}L(\mathcal{F}_{s}(i^{nes}),e(\pi^{j})))}{||\nabla_{i^{\prime}}L(\mathcal{F}_{s}(i^{nes}),e(\pi^{j}))||_{1}}\\ i^{j+1}=\mathrm{Clip}^{\epsilon}_{i}\{i^{j}+\alpha_{i^{\prime}}\cdot\mathrm{sign}(g^{j+1}_{i^{\prime}})\}\\ \end{array}\right.$ (11)
where $T$ is the stochastic transformation function with the probability $p$, $\alpha_{i^{\prime}}$ is the step size, $\mu$ is the decay factor of the momentum term, $g^{j}_{i^{\prime}}$ is the accumulated gradient at iteration $j$, and $\mathrm{Clip}^{\epsilon}_{i}$ indicates the resulting adversarial images are clipped within the $\epsilon-$ball of the original image $i$.
@@ IV-A Experimental Settings
Evaluation Metrics. For evaluation, we report (1) the attack success rate (ASR), which measures the rate of successful adversarial samples. Similar to Co-attack [17], we define the ASR as the results when the original samples are fed into the target models minus the results when the adversarial samples are fed into the target models. (2) Semantic Similarity (Sim.), which is computed between the original and adversarial sentences and is commonly approximated by the universal sentence encoder [44]. (3) Tokens error rate (TER.), which is the proportion of words or characters in a text that have been altered to create an adversarial text, generally, less perturbation results in more semantic consistency. Note: Semantic Similarity (Sim.) and Token Error Rate (TER) are used to evaluate the extent of change between original and adversarial texts, identifying methods with superior ASR in a black-box setting under equivalent semantic deviation and token errors. The whole method is implemented by Pytorch [45] and all experiments are conducted in a GeForce RTX A6000 GPU.